To four significant figures, a) what is the energy of a photon emitted from a hydrogen atom when an electron makes a transition from n=6 to n=2

b) find the energy of electromagnetic radiation necessary for ionization (transition to n=invinity) of the 1s1 electron in a hydrogen atom?

Please help step by step so I can duplicate it.

Thank you
Answer #1

a.

Rydberg equation is

= RH × ( )

RH is Rydberg constant. h is planck constant , c is velocity of light.

ni =2

nf = 6

or, E = R​​​​​​H × h× c× ()

or, E = 1.09737 × 107 × 6.626×10-34 ×3×108 ×(2/9)

or, E = 4.847 ×10-19 J.

hence, energy of photon emitted = 4.847×10-19 J.

b.

ni = 1

nf = infinity

E = RH × C× h × ( )

or, E = 1.09737×107 ×6.626×10-34 × 3×108 ×1

or, E = 2.181×10-18 J.